# Fully connected twin-field quantum key distribution network

### v1

@@ Abstract
Quantum key distribution (QKD) is a key application in quantum communication, enabling secure key exchange between parties using quantum states. Twin-field (TF) QKD offers a promising solution that surpasses the repeaterless limits, and its measurement-device-independent nature makes it suitable for star-type network architectures. In this work, we propose and demonstrate a fully connected TF-QKD network architecture, where users prepare quantum signals and send them to network nodes. These nodes use splitters to route the signals to measurement units, enabling secure key distribution among arbitrary users. A proof-of-principle demonstration with three users successfully achieved secure key sharing over simulated link losses of up to 30 dB, with an average rate of $5.01\times 10^{-7}$ bit per pulse. Additionally, simulations show that the architecture can support up to $32$ users with a secure key rate of 280.90 bits/s under typical fiber loss conditions. This approach represents a significant advancement in the topology of untrusted-node QKD networks and holds promise for practical, large-scale applications in secure communication.
@@ I Introduction
Quantum key distribution (QKD) is one of the most mature applications of quantum information processing [1]. It uses quantum states to securely distribute symmetric keys between communication parties. Since the first QKD protocol was proposed [2], significant efforts have steadily advanced QKD technology [3, 4, 5, 6, 7, 8]. Recently, a record-breaking achievement was made, with secure key rates exceeding tens of megahertz [5, 6]. Trusted-node QKD networks [9, 10, 11, 12, 13, 14], including integrated space-ground quantum communication networks [14], have been rapidly advancing. However, challenges persist, particularly in long-distance transmission and network expansion, both of which are crucial for the wider deployment of QKD.
Figure 1: (a) Schematic of the physical layer of the fully connected TF QKD network architecture for $N$ users. (b) Schematic of the user transmitter device, which includes frequency-stabilized coherent light based on a local frequency references, intensity modulations, and phase modulations. (c) Schematic of the network node configuration, which includes the same number of polarization alignment modules and $1\times(N-1)$ polarization-maintaining splitters, along with an optical switch array and measurement units satisfying a fully connected setup.
@@
In this study, we propose and demonstrate a fully connected TF-QKD network architecture. Our architecture is inspired by recent advancements in protocols [36, 37, 38] and open-architecture TF-QKD schemes [39, 40, 41, 42, 43], eliminating the need for additional channels and devices to synchronize the frequencies of independent light sources, thereby offering greater flexibility and scalability. Furthermore, each user prepares quantum signals and sends them to a network node in single-path transmission, which inherently provides immunity to bidirectional transmission security vulnerabilities. Furthermore, the node employs beam splitters to route the received quantum signals to different single-photon interference measurement units, enabling secure key distribution among arbitrary users. We conduct a proof-of-principle demonstration of a three-user fully connected TF-QKD network, where secure key sharing is successfully achieved between any pair of users over a simulated link loss of up to $30$ dB, with an average rate of $5.01\times 10^{-7}$ bit per pulse. Additionally, we perform simulations based on the parameters of the experimental system. The results show that our network architecture can support up to $32$ users in a symmetric scenario, with a secure key rate of $280.90$ bits/s between user pairs under a fiber channel loss of $20$ dB. This network architecture enhances the topology of untrusted-node QKD networks and promotes their practical applications.
@@ II Network architecture
The fully connected TF QKD network architecture is depicted in Figure 1. The physical layer adopts a star topology consisting of end users, fiber channels, and untrusted network node, as shown in Figure 1(a). Each user is equipped with a transmitter, located at the terminal of channel extending from the central node, as shown in Figure 1(b). The network node includes polarization monitors, polarization-maintaining splitters, an optical switch array, and single-photon interference measurement units. Full connectivity among multiple users is achieved through the designed of nodal components, as exemplified in Figure 1(c).
@@
Figure 2: Schematic of our experimental setup. A commercial laser serves as the light source for all users. Each frame of light pulse is modulated by the first intensity modulation module (IM-$1$) into $400$ signal pulses, $600$ reference pulses, and $24$ vacuum pulses. The intensity modulation module includes beam splitter (BS) and a phase modulator (PM). Subsequently, each user employs a second intensity modulation module (IM-$2$) and a phase modulator (PM) to implement encoding, decoy intensity modulation, and discrete random phase modulation on the optical pulses. The pulses are then attenuated via a variable optical attenuator (VOA), followed by a simulated channel using VOA. At the network node, quantum signals from user pass through polarization controller (PC) and polarization beam splitter (PBS), then couple into a polarization-maintaining splitter (PMS). The signals from the user pairs interfere at a BS within the measurement unit (MU). Finally, the quantum signals are detected by superconducting nanowire single-photon detectors (SNSPDs).
An $N$-user network instance is considered here to further elucidate our network. Each user’s transmitter comprises a laser source, intensity modulators, and phase modulators. The user prepares frequency-stabilized light signal based on a local frequency reference, then modulate the light signal’s intensity and phase according to the specific TF-QKD protocol. The modulated signal is transmitted to network nodes via fiber channel.
The polarization monitor at the network node tracks and compensates for polarization drift in the user signal. The signal is then aligned and coupled into a $1\times(N-1)$ polarization-maintaining splitter, where it is randomly directed to one of its branches. The $N-1$ branches of the splitter enable connections between the user and the remaining $N-1$ users. The $(N^{2}-N)\times(N^{2}-N)$ optical switch array allocates input ports to the branches of all splitters. For example, any user who intends to communicate with User-11{1}1 can establish a complete quantum channel by accessing the measurement unit through the optical switch. For an $N$-user network, up to $N\times(N-1)/2$ measurement units are required to achieve full connectivity. Figure 1(c) shows an $N$-user fully connected configuration. Considering some dedicated line communication scenarios, network node can also form a static configuration by directly wiring splitters and measurement units, as demonstrated in the three-user network in Section III.
@@
In this network, key distribution between any pair of users is implemented based on the TF-QKD protocol, which requires coherent control of the light fields from remote users. The frequency difference between independent user lasers or phase shifts induced by channel fluctuations affect the coherence of these light fields. Building on recent advancements [43], users can employ acetylene-stabilized laser, eliminating the need for a shared optical frequency reference. This makes it possible for the independent lasers of users in the network to maintain frequency alignment using a local frequency reference. Furthermore, there are mature solutions to effectively compensate for the phase shifts caused by fiber fluctuations over several hundred kilometers, either through real-time feedback [26, 27, 39] or post-processing techniques [28, 43].
@@ III Experiment
To validate the proposed architecture, we build a proof-of-principle experimental setup, which includes three fully connected users in a TF-QKD network. Figure 2 shows the schematic of experimental setup. A laser source operating at $1550$ nm generates optical pulses at a $100$ MHz clock rate, with a pulse width of approximately $500$ ps. These pulses enter the first intensity modulation module (IM-$1$), which is a Sagnac-based interferometer. The optical pulses are modulated into sequential frame-based signals. Specifically, each frame consists of $1024$ optical pulses, comprising $400$ signal pulses, $600$ reference pulses and $24$ vacuum pulses. These pulses are then sent to the three users.
Each user encodes the signal pulses according to the three-intensity sending-or-not-sending (SNS) TF QKD protocol [44, 45] by employing the second intensity modulation module (IM-$2$), thereby producing a signal window for key extraction and a decoy window to estimate information leakage. Subsequently, a phase modulator (PM) is employed to perform $16$-level random phase modulation over a range of $2\pi$ to the signal pulses. Finally, the signal pulses are attenuated to the single-photon level using a variable optical attenuator (VOA) before entering the channel.
Figure 3: Interference results between users in a three-user fully connected network within $200$ ms. The interference results are captured by recording the counts of the detectors D0 and D1. The average interference visibilities for user pairs $1$-$2$, $1$-$3$, and $2$-$3$ are $90.87\%$, $95.82\%$, and $96.72\%$, respectively.
Signals from different users reach the network node through a channel simulated by a VOA. The arriving quantum signals are first processed by a polarization controller (PC) to align their polarization to the one of orthogonal modes of the followed polarization beam splitter (PBS). The signals are then efficiently coupled into a $1\times 2$ fast-axis block polarization-maintaining splitter (PMS), and randomly routed to different branches with the same polarization. The signals from the user pairs interfere with the same polarization at the beam splitter (BS) within the measurement unit (MU). The interference outputs between user pairs are detected by two superconducting nanowire single-photon detectors (SNSPDs), with detection efficiency around $45\%$ and a dark count rate of about $8\times 10^{-8}$. The detected photon events are recorded by the TDC and then processed by a computer. The detection events of the reference pulse are used in post-processing to estimate the phase fluctuation between users. Within the effective arrival time $t$, the detection events in the decoy window are further filtered according to the following condition:
@@ IV Results
Using the described network system, we first measure the interference between users to preliminarily assess system performance. The measurement is conducted with the phase modulator installed but not actively modulating. The interference results are obtained by monitoring the counts from detectors in the MU, as shown in Figure 3. We observe that the average interference visibilities for user pairs $1$-$2$, $1$-$3$, and $2$-$3$ are $90.87\%$, $95.82\%$, and $96.72\%$, respectively, over a period of approximately $200$ ms. Due to the single laser scheme adopted, phase fluctuations affecting interference visibility in the experiment primarily originate from optical components.
@@
Figure 4: (a) Experimental results in a three-user fully connected network. (b) The secure key rate per user pair in a $32$-user fully connected network.
@@
To evaluate the network’s performance with an increasing number of users, we consider our system parameters along with the splitter loss and clock rate parameters from reference [46]. The network performance in a symmetric scenario is shown in Figure 4(b). The results indicate that this network architecture can support a $32$-user fully connected network, achieving secure key rates of $490.63$ bits/s per user pair for a simulated fiber loss of $10$ dB and $208.90$ bits/s per user pair for a simulated fiber loss of $20$ dB.
@@ V Conclusion
In summary, we propose a fully connected TF-QKD network architecture, designed to efficiently support secure key distribution among multiple users without relying on trusted nodes. Each user is connected to a central node equipped with splitters, which randomly route quantum signals to measurement units, enabling secure key sessions between arbitrary user pairs. A three-user experimental demonstrate the feasibility of this network expansion strategy, achieving stable secure key distribution across all pairwise connections. Furthermore, simulations show that this network can support up to $32$ users with a secure key rate of $280.90$ bits/s for user pairs under a simulated channel loss of $20$ dB.
@@
Acknowledgment
This study was supported by the National Natural Science Foundation of China (Nos. 62171144 and 62031024), Guangxi Science Foundation (Nos. 2025GXNSFAA069137 and GXR-1BGQ2424005), and Innovation Project of Guangxi Graduate Education (No. YCSW2022040).

### v2

# Scalable twin-field quantum key distribution network enabled by adaptable architecture

@@ Abstract
Quantum key distribution (QKD) is a key application in quantum communication, enabling secure key exchange between parties using quantum states. Twin-field (TF) QKD offers a promising solution that surpasses the repeaterless limits, and its measurement-device-independent nature makes it suitable for star-type network architectures. In this work, we propose a scalable TF-QKD network with adaptable architecture, where users prepare quantum signals and send them to network nodes. These nodes use an optical switch to route the signals to multi-user measurement units, enabling secure key distribution among arbitrary users and adapting to complex connection demands of the network. A proof-of-principle demonstration with three users successfully achieved secure key sharing over simulated link losses of up to $30$ dB, with an average rate of $19.57$ bit/s. Additionally, simulations show that the proposed architecture can achieve a total secure key rate of $4.84\times 10^{4}$ bit/s at $100$ km in a symmetric $32$-user network. This approach represents a significant advancement in the topology of untrusted-node QKD networks and holds promise for practical, large-scale applications in secure communication.
@@ I Introduction
Quantum key distribution (QKD) is one of the most mature applications of quantum information processing [1]. It uses quantum states to securely distribute symmetric keys between communication parties. Since the first QKD protocol was proposed [2], significant efforts have steadily advanced QKD technology [3, 4, 5, 6, 7, 8, 9]. Recently, a record-breaking achievement was made, with secure key rates exceeding tens of megahertz [5, 6]. Trusted-node QKD networks [10, 11, 12, 13, 14, 15], including integrated space-ground quantum communication networks [15], have been rapidly advancing. However, challenges persist, particularly in long-distance transmission and network expansion, both of which are crucial for the wider deployment of QKD.
Figure 1: (a) Schematic of the scalable TF-QKD network architecture for $N$ users. The physical layer adopts a star topology consisting of end users, quantum channels, and untrusted network node. Each user requires only a transmitter. The network node utilizes an optical switch to manage user connections, directing quantum signals from users to designated measurement units (MUs). (b) Schematic of $n$-user MU with two extended ports per user output. (c) Schematic of $n$-user MU with three extended ports per user output.
@@
In this study, we propose a scalable TF-QKD network with adaptable architecture. Our architecture is inspired by recent advancements in protocols [37, 38, 39] and open-architecture TF-QKD schemes [40, 41, 42, 43, 44], eliminating the need for additional channels and devices to synchronize the frequencies of independent light sources, offering greater flexibility and scalability. Furthermore, each user prepares quantum signals and sends them to a network node in single-path transmission, which inherently provides immunity to bidirectional transmission security vulnerabilities. Furthermore, the network node employs an optical switch to dynamically group users and, in conjunction with multi-user measurement units (MUs), flexibly supports secure key distribution for the network. Within multi-user MU, the output port of each user is expanded using a multi-port splitter, enabling simultaneous key generation among a larger number of user pairs.
We conduct a proof-of-principle demonstration of a three-user TF-QKD network, in which secure key exchange is successfully achieved between any pair of users over simulated link losses of up to $30$ dB, with an average rate of $5.01\times 10^{-7}$ bit/pulse or $19.57$ bit/s. Additionally, we perform simulations based on the parameters of the experimental system. The results show that the proposed architecture can achieve a total secure key rate of $1.24\times 10^{-3}$ bit/pulse or $4.84\times 10^{4}$ bit/s at $100$ km in a symmetric $32$-user network with a specific MU configuration. This network architecture enhances the topology of untrusted-node QKD networks and promotes their practical applications.
@@ II Network architecture
Figure 1(a) depicts the adaptive architecture of the TF-QKD network. The network adopts a star topology centered on the network node, with users located at the terminals of quantum channels extending from the node. Each user connects to the TF-QKD network as a transmitter that prepares quantum state according to specific TF-QKD protocols. The network node contains an $N\times N$ optical switch and multiple MUs. The optical switch groups arbitrary users and routes their quantum signals to the same MU, such as the $2$-user and $3$-user MUs. Figure 1(b) and (c) show $n$-user MU configurations featuring varying numbers of extended ports per user output. The MU expands each user’s output to support their connections with multiple users. The specific number of extended ports depends on user’s demands.
@@
Figure 2: Schematic of our experimental setup. A commercial laser serves as the light source for all users. Each frame of light pulse is modulated by the first intensity modulation module (IM-$1$) into $400$ signal pulses, $600$ reference pulses, and $24$ vacuum pulses. The intensity modulation module includes beam splitter (BS) and a phase modulator (PM). Subsequently, each user employs a second intensity modulation module (IM-$2$) and a phase modulator (PM) to implement encoding, decoy intensity modulation, and discrete random phase modulation on the optical pulses. The pulses are then attenuated via a variable optical attenuator (VOA), followed by a simulated channel using VOA. At the network node, quantum signals from user pass through polarization controller (PC) and polarization beam splitter (PBS), then couple into a polarization-maintaining splitter (PMS). The signals from the user pairs interfere at a BS within the single-photon Bell state measurement (S-BSM) module. Finally, the quantum signals are detected by superconducting nanowire single-photon detectors (SNSPDs).
@@
In this network, key distribution between any pair of users is implemented based on the TF-QKD protocol, which requires coherent control of the light fields from remote users. The frequency difference between independent user lasers or phase shifts induced by channel fluctuations affect the coherence of these light fields. Building on recent advancements [42], users can employ commercial kilohertz-linewidth semiconductor lasers to successfully perform TF-QKD. This scheme reconstructs a global phase reference through post-processing, eliminating the need for additional service fibers for frequency dissemination and thereby significantly simplifying hardware requirements and network deployments. Accordingly, the transmitter typically includes a commercial laser, multiple intensity modulators, and phase modulators. Furthermore, there are mature solutions to effectively compensate for the phase shifts caused by fiber fluctuations over several hundred kilometers, either through real-time feedback [27, 28, 40] or post-processing techniques [29, 42, 44].
The network node employs an optical switch and multiple MUs to support any-to-any user connectivity and adapt to complex, real-world network topologies. Specifically, the reconfigurable optical switch enables adaptable grouping of users through multiple MUs, subject to the constraint $\sum_{2}^{n}nM_{n}<N$, where $N$ denotes the number of optical switch output ports, $n$ represents the output ports occupied by an $n$-user MU, and $M_{n}$ indicates the quantity of $n$-user MUs deployed.
For instance, a $2$-user MU essentially corresponds to the single-photon Bell state measurement (S-BSM) module used in typically TF-QKD system. A $3$-user MU extends each user’s output to two ports, enabling simultaneous pairwise connectivity among all users. For a $n$-user MU, as shown in Fig. 1(b), each user’s output is extended to two ports, allowing the formation of up to $n$ distinct user pairs within the unit. Under this modular MU configuration, the total number of supported concurrent user pairs becomes $M_{2}+\sum_{3}^{n}nM_{n,2}$, where $M_{n,2}$ denotes the number of $n$-user MUs with two extended ports per user. Intuitively, increasing the number of output ports per user enables more user pairs to be supported simultaneously, enhancing adaptability to diverse connection demands in real-world scenarios. As shown in Fig. 1(c), an $n$-user MU with three output ports per user can support up to $\lfloor 3n/2\rfloor$ pairs of users. Incorporating such MUs into the network increases the maximum number of concurrent connections to $M_{2}+\sum_{3}^{n}nM_{n,2}+\sum_{4}^{n}\lfloor 3n/2\rfloor M_{n,3}$, where $M_{n,3}$ is the number of $n$-user MUs with three extended ports per user.
More generally, when MUs with varying numbers of extended ports are integrated, the network’s maximum concurrent connections $C$ can expressed as:
$C=M_{2}+\sum_{3}^{n}nM_{n,2}+\sum_{4}^{n}\sum_{i=3}^{n-1}\lfloor ni/2\rfloor M_{n,i},$ (1)
subject to the resource constraint $2M_{2}+\sum_{3}^{n}n\sum_{i=2}^{n-1}M_{n,i}<N$. Here, $M_{n,i}$ denotes the number of $n$-user MUs with $i$ extended ports per user. These modular MU configurations, when integrated with a reconfigurable optical switch, significantly enhance the network’s scalability and flexibility. They enable the system to dynamically allocate resources and satisfy varying user connection demands while adhering to the optical switch’s total output port constraints.
@@ III Experiment
Figure 3: Interference results between users in a three-user network within $200$ ms. The interference results are captured by recording the counts of the detectors D0 and D1. The average interference visibilities for user pairs $1$-$2$, $1$-$3$, and $2$-$3$ are $90.87\%$, $95.82\%$, and $96.72\%$, respectively.
We build a proof-of-principle experimental setup in which three users are statically connected to a three-user MU, employing a single-laser scheme rather than fully independent lasers. Despite this simplification, our experiment can serve as a functional instantiation of the adaptive network architecture components. Figure 2 shows the schematic of experimental setup. A laser source operating at $1550$ nm generates optical pulses at a $100$ MHz clock rate, with a pulse width of approximately $500$ ps. These pulses enter the first intensity modulation module (IM-$1$), which is a Sagnac-based interferometer. The optical pulses are modulated into sequential frame-based signals. Specifically, each frame consists of $1024$ optical pulses, comprising $400$ signal pulses, $600$ reference pulses and $24$ vacuum pulses. These pulses are then sent to the three users.
Each user encodes the signal pulses according to the three-intensity sending-or-not-sending (SNS) TF QKD protocol [22, 45, 46] by employing the second intensity modulation module (IM-$2$), thereby producing a signal window for key extraction and a decoy window to estimate information leakage. Subsequently, a phase modulator (PM) is employed to perform $16$-level random phase modulation over a range of $2\pi$ to the signal pulses. Finally, the signal pulses are attenuated to the single-photon level using a variable optical attenuator (VOA) before entering the channel.
Figure 4: (a) Experimental results in a three-user network. (b) The total secure key rate of the network over varying transmission distances and different numbers of active users.
Signals from different users reach the three-user MU of network node through a channel simulated by a VOA. The arriving quantum signals are first processed by a polarization controller (PC) to align their polarization to the one of orthogonal modes of the followed polarization beam splitter (PBS). The signals are then efficiently coupled into a $1\times 2$ fast-axis block polarization-maintaining splitter (PMS), and randomly routed to different branches with the same polarization. The signals from the user pairs interfere with the same polarization at the beam splitter (BS) within the S-BSM module. The interference outputs between user pairs are detected by two superconducting nanowire single-photon detectors (SNSPDs), with detection efficiency around $45\%$ and a dark count rate of about $8\times 10^{-8}$. The detected photon events are recorded by the TDC and then processed by a computer. The detection events of the reference pulse are used in post-processing to estimate the phase fluctuation between users. Within the effective arrival time $t$, the detection events in the decoy window are further filtered according to the following condition:
@@ IV Results
Using the described experimental system, we first measure the interference between users to preliminarily assess system performance. The measurement is conducted with the phase modulator installed but not actively modulating. The interference results are obtained by monitoring the counts from detectors in the MU, as shown in Figure 3. We observe that the average interference visibilities for user pairs $1$-$2$, $1$-$3$, and $2$-$3$ are $90.87\%$, $95.82\%$, and $96.72\%$, respectively, over a period of approximately $200$ ms. Due to the single-laser scheme adopted, phase fluctuations affecting interference visibility in the experiment primarily originate from optical components.
@@
To evaluate the network’s performance with a larger number of users, we conduct numerical simulation of a network that can support up to $32$ users in a symmetric scenario. We consider our system parameters along with the loss value of $1\times 8$ splitter from reference [47]. The simulation assumes a specific MU configuration of $\{M_{2}=1,M_{3,2}=1,M_{4,2}=1,M_{6,2}=1,M_{8,2}=1,M_{9,8}=1\}$ and a total of $N=10^{11}$ pulses sent by each user. The total secure key rate of the network is evaluated with respect to transmission distance and the number of active users, as depicted in Fig. 4(b). As the number of active users increases, we determine the maximum number of user pairs that can be simultaneously supported under the given MU configuration, and calculate the corresponding total secure key rate of the network.
The independent S-BSMs within the MUs at the network node improve the overall network’s capacity for parallel key generation. The fully loaded network achieves a total secure key rate of $1.24\times 10^{-3}$ bit/pulse or $4.84\times 10^{4}$ bit/s at a transmission distance of $100$ km. The achievable secure key rate between user pairs can reach up to $4.77\times 10^{3}$ bit/s, which meets the bit rate requirement for real-time one-time-pad encryption of voice communications. Even for the worst-performing user pairs, a secure key rate of approximately $256.86$ bit/s can be achieved under the quantum transmission duty cycle used in our experiment.
With nine active users, the $9$-user MU with eight extended ports enables up to $28$ pairs of users simultaneously. Although this MU allows for a greater number of concurrent connections, it also requires significantly more detection resources. Moreover, due to the higher loss in multi-port MUs, the key rate of each user pair becomes more susceptible to statistical fluctuations at the same transmission distance, thereby limiting the network performance, as shown in Fig. 4(b). Therefore, it is crucial to optimize the MU composition based on practical resource constraints and user demands to achieve a balance between flexibility, resource usage, and network performance.
@@ V Conclusion
In summary, we propose a scalable TF-QKD network that supports adaptive user pairing based on a modular design. By integrating an $N\times N$ optical switch with multi-user MUs of varying port configurations, our network supports flexible connectivity and concurrent key generation among multiple users. We perform a proof-of-principle network experiment with a three-user MU and demonstrate secure key distribution between any pairs of users over simulated fiber losses up to $30$ dB. Furthermore, numerical simulation shows that the proposed architecture can achieve a total secure key rate of $4.84\times 10^{4}$ bit/s over a distance of $100$ km in a $32$-user network.
@@
Acknowledgment
This study was supported by the National Natural Science Foundation of China (Nos. 62171144 and 62031024), Guangxi Science Foundation (Nos. 2025GXNSFAA069137 and GXR-1BGQ2424005), and Innovation Project of Guangxi Graduate Education (No. YCBZ2025064).